# Divergences in the hadronic light-by-light amplitude of the holographic soft-wall model

### v1

@@ II Pseudoscalar and axial sector in the soft-wall model
for all $z$ without backreactions to the metric, as this is the most commonly used form in the literature. After understanding the calculations with this background, the analysis can be readily generalized to more general $X_{0}$, warp factor $w(z)$, and dilaton profile, which will be analyzed in section VI. As will be seen there, when $X_{0}\rightarrow\text{const.}$ at infinity, there will be even problems with the definition of individual transition form factors. Restricting ourselves to the $N_{f}=2$ case, the parameters $c,m_{q}=m_{u,d},\sigma$ are fitted to the $\rho$-meson mass, the $\pi^{0}$ mass and decay constant, respectively, while $g_{5}=2\pi$ follows from a fit of the vector-current correlator to the OPE of QCD.
@@
$\partial_{z}(\frac{e^{-\phi}}{z}\partial_{z}A^{\perp}_{\mu})+\frac{m_{n}^{2}}{z}e^{-\phi}A^{\perp}_{\mu}-g_{5}^{2}\frac{v^{2}}{z^{3}}e^{-\phi}A^{\perp}_{\mu}=0.$ (8)
@@
which also leads to a Schrödinger equation, now with a potential
@@
$\tilde{V}_{\mu}(z,q)=\mathcal{J}(z,q^{2})P_{\mu}^{\;\nu}(q)V^{(0)}_{\nu}(q)+q_{\mu}\frac{q^{\nu}V^{(0)}_{\nu}(q)}{q^{2}},$ (16)
with $V^{(0)}$ being the boundary value and $P^{\mu\nu}(q)=\eta^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}$. The bulk-to-boundary propagator $\mathcal{J}(z,q^{2})$ obeys the equation
$\partial_{z}\left(\frac{e^{-\phi}}{z}\partial_{z}\mathcal{J}(z,q^{2})\right)+\frac{q^{2}e^{-\phi}}{z}\mathcal{J}(z,q^{2})=0,$ (17)
and $\mathcal{J}(0,q^{2})=1$. The two-point function is given by
$\Pi_{ab}^{\mu\nu}(q)=\frac{1}{g_{5}^{2}}\delta_{ab}P^{\mu\nu}(q)\frac{\partial_{z}\mathcal{J}(z,q^{2})}{z}\bigg|_{z\to 0}.$ (18)
In order to avoid a pole at $q^{2}=0$ of the form $\Pi^{\mu\nu}_{ab}\sim\delta_{ab}q^{\mu}q^{\nu}/q^{2}$ from massless scalars in $V_{\mu}\propto\partial_{\mu}S$, one needs $z^{-1}{\partial_{z}\mathcal{J}(z,0)}|_{z\to 0}$ to vanish [^2] The familiar divergence ${\partial_{z}\mathcal{J}(\varepsilon,q^{2})}/{\varepsilon}\sim q^{2}\log(\varepsilon Q)$, which is canceled by holographic renormalization, is not relevant in this discussion as it only appears at nonzero $q^{2}$. . For $a<0$, the most general regular solution of the EOM obeying the UV b.c. at $q^{2}=0$ is
@@
For $a>0$, the second term needs to vanish since it would grow at large $z$, but for $a<0$, any $c_{2}$ seems to be allowed. In order to determine $c_{2}$, we look at small but nonzero values of $q^{2}$ and demand that $\mathcal{J}(z,q^{2})$ is regular as $q^{2}\rightarrow 0$. It is then immediately seen that when expanding $\mathcal{J}(z,q^{2})=\mathcal{J}(z,0)+q^{2}\mathcal{J}^{(1)}(z,q^{2})+$... the perturbation $\mathcal{J}^{(1)}$ grows logarithmically as $z\rightarrow 0$ if $1+c_{2}\int_{0}^{\infty}dz\,z\,e^{\phi}\neq 0$. Demanding regularity means that
@@ III.1 WKB and definition of four regions
However, even at very large $E_{n}$, there will always be regions where $\frac{k^{\prime}}{k^{2}}\ll 1$ is not obeyed, and we need to find different ways to approximate the solution. These regions are the small-$z$ region, where $V_{P,A}(z)\gg E_{n}$ and therefore $\frac{k^{\prime}}{k^{2}}\sim\frac{1}{z^{4}}$ (the size of this region shrinks, as $E_{n}$ grows), and the regions around the points $z_{l},z_{r}$ where $k^{2}(z)=0$. The points $z_{l},z_{r}$ are approximately given by
@@
In appendix A we have spelled out explicitly also the analysis for the axial vector mesons, which differs from the pseudoscalar case in a rather minimal way. The potential $V_{A}$ is slightly different at small $z$, leading also to a mass spectrum $M^{A}\sim n^{2/3}$, whereas the normalization condition (15) does not involve $E_{n}=(M^{A}_{n})^{2}$.
@@ III.2 Transition form factors, sum rules and hadronic light-by-light tensor
Even without the use of the WKB approximation, one can see signs of this pathological behavior in the g2g-2 contribution by simply computing its value for a sufficiently large number of excited states numerically. The results (with $m_{q}$ such that $m_{\pi^{0}}=135$ MeV) are shown in Table 1 where one can clearly see an increase in the $a_{\mu}$ contributions from excited pseudoscalars with masses above $7.5$ GeV, preventing a convergence of their sum.
@@
The analysis for large $m_{n}$ is very similar to the pseudoscalar case one, so we will not repeat it in detail here. Going through the steps as in the pseudoscalar case, we find the scaling
@@
The WKB machinery may also be used for hard-wall models to compute transition form factors for large $n$. We give closed formulas for $A_{n}(Q^{2},Q^{2})$ and $A_{n}(Q^{2},0)$ in B and comment on the realization of the Melnikov-Vainshtein constraint. The asymptotic scaling laws of [54] are recovered for $Q\gg\sqrt{E_{n}}=m_{n}$, but in this limit they do not contribute to the short-distance constraints anymore.
@@ IV VVA correlator and non-commutativity of limits
$\times\bigg(q_{2}^{\alpha}\mathcal{J}^{\prime}(z,q_{1}^{2})\mathcal{J}(z,q_{2}^{2})-q_{1}^{\alpha}\mathcal{J}^{\prime}(z,q_{2}^{2})\mathcal{J}(z,q_{1}^{2})\bigg).$ (66)
$\mathcal{A}^{\perp}(z,q)$ solves the axial equations of motion (8) with $m_{n}^{2}$ replaced by $q^{2}$, subject to the boundary condition $\mathcal{A}^{\perp}(0,q)=1$. Similarly $\varphi(z,q)$ together with $\eta(z,q)$ solve the pseudoscalar equations (6) (with $m_{n}^{2}\rightarrow q^{2}$) obeying $\varphi(0,q)=1$, $\eta(0,q)=0$.
@@
$\bar{\mathcal{W}}_{2}^{\mathrm{pole}}(Q^{2})$ $=-\frac{N_{c}}{2\pi^{2}}\text{tr}\left(\mathcal{Q}^{2}\mathcal{Q}_{5}\right)\frac{F_{A}}{Q^{2}+m_{A}^{2}}\frac{A(Q^{2},0)}{2}$
$=-\frac{N_{c}}{2\pi^{2}}\text{tr}\left(\mathcal{Q}^{2}\mathcal{Q}_{5}\right)\frac{F_{A}}{Q^{2}+m_{A}^{2}}$
@@
with $F^{A}$ being an axial-vector decay constant and $A(Q^{2},0)$ the singly virtual axial TFF. At large $N_{c}$, the only singularities one would expect in correlation functions and amplitudes (at leading order) are poles, and a dispersive analysis would reconstruct the VVA correlator simply as a sum over all such pole contributions (IV) (except for potential subtractions). According to the discussion in the previous section, the TFF $A_{n}(Q^{2},0)$ scales exponentially, $\sim e^{\frac{c^{2}}{2g_{5}\sigma}\sqrt{E}_{n}}$, and the sum over all pole contributions does not converge. This is seemingly in contradiction to the above analysis, where it was concluded that $\bar{\mathcal{W}}_{2}$ is finite when calculated using the axial bulk-to-boundary propagator. In Appendix D, we trace this back to a lack of uniform convergence of the mode expansion, which does not permit to swap the infinite sum over modes with the indefinite integration over the holographic coordinate $z$.
@@ V Hadronic light-by-light tensor from the 5-dimensional approach
Noting that the Wronskian $\xi_{L}^{\prime}(z^{\prime})\xi_{R}(z^{\prime})-\xi_{L}(z^{\prime})\xi_{R}^{\prime}(z^{\prime})$ is independent of $z^{\prime}$, the symmetry $G_{A}(z,z^{\prime},Q)=G_{A}(z^{\prime},z,Q)$ is apparent. Explicit formulas for $\xi_{L},\,\xi_{R}$ in the WKB approach are given in Appendix C.
@@
By symmetry, we only have to integrate below the diagonal [^8] The other $z,\,z^{\prime}$ dependent factors from the vector bulk-to-boundary propagators are not symmetric in $z,\,z^{\prime}$ for generic $Q_{i}$, but inside the integral only the symmetric part contributes.. Below the diagonal at large $z,z^{\prime}$, the WKB approximation detailed in Appendix C predicts the axial bulk to bulk propagator to scale like
@@ VI.1 Background modifications
for fixed [^11] We have neither determined the power-law prefactor $m_{n}^{\lambda}$ nor the $Q_{1},Q_{2}$ dependence, as the exponential already proves the divergence of the HLbL tensor. $Q_{1},Q_{2}$ by similar arguments as in section III.2.
If the second term in (95) contributes at the same order or even dominates, the individual TFFs will also be well defined, but exponential scaling similar to (1) will still occur. Also the axial-vector spectrum will not lie on a linear Regge trajectory (except when $w(z^{2})v(z)^{2}\sim z^{2}$, but then with a different slope than vector mesons). The exponent in the scaling law will in general also not be proportional to $m_{n}$ anymore, rather if the second term in the potential goes like $z^{\alpha}$, with $\alpha\geq 2$, the exponent will be proportional to $m_{n}^{4/\alpha}$. As $\alpha$ gets larger, the factor $e^{m_{n}^{4/\alpha}}$ will get smaller at fixed $m_{n}$, but as $m_{n}\rightarrow\infty$ it will still dominate any power-law factor and observables such as the light-by-light tensor will diverge. In the case that the $w^{2}(z)v^{2}(z)$ factor behaves exponentially at large $z$, the point $z_{r}$ will scale as $z_{r}\sim\left(\log(E_{n})\right)^{1/b}$, for some $b>0$. In that situation, the scaling of the TFF with $E_{n}$ becomes a power law, and a more sophisticated analysis is needed to decide whether the infinite sums discussed in this paper converge. It seems plausible that there are ranges of $b$ for which convergence can be achieved. Another interesting case is when $w^{2}(z)v^{2}(z)$ has a divergence at some finite $z=z_{0}$. In that case, $z_{r}$ will be equal to $z_{0}$ for all $n$ larger than some critical $n_{c}$. The exponential factor $e^{\frac{c^{2}}{2}z_{r}^{2}}$ is then completely harmless and the spectrum becomes hard-wall like [^12] To have well defined individual TFFs one additionally needs $w(z)^{2}v^{2}(z)\sim az^{\beta}$ for $z>z_{0}$ with $a>0,\beta\geq 2$. . As briefly discussed at the end of the following subsection, the relevant sums all converge for hard-wall models.
@@ VI.2 Scalar-extended CS term
Hard-wall models do not suffer from the above exponential scaling of TFFs since the modes are restricted to a finite region $z\in[0,z_{0}]$. [^14] This is also the case for the Witten-Sakai-Sugimoto model when brought into Schrödinger form. There, however, the geometry is not asymptotically AdS so that one cannot match to the short-distance constraints of QCD. This leads to the scaling behaviors $F(0,0)\sim E_{n}^{-1},f_{n}\sim E_{n}^{-1/4}$ with $E_{n}\sim n^{2}$, which leads to a convergent sum rule (53), and a convergent light-by-light tensor for the pseudoscalars (the axials are analogous). Moreover, the experimental data for the pion TFF (and also for the $a_{1}$) turn out to be reproduced surprisingly well [18] with a standard Chern-Simons term. [^15] In the HW model, a scalar-extended Chern-Simons term with $\alpha^{\prime}=\frac{1}{6}$ helps to improve some features in the $f_{1}$-$f_{1}^{\prime}$ sector, but also leads to an unsatisfactory pion TFF [20].

### v2

@@ II Pseudoscalar and axial sector in the soft-wall model
for all $z$ without backreactions to the metric, as this is the most commonly used form in the literature. After understanding the calculations with this background, the analysis can be readily generalized to more general $X_{0}$, warp factor $w(z)$, and dilaton profile, which will be analyzed in section VI. As will be seen there, when $X_{0}\rightarrow\text{const.}$ at infinity, there will be even problems with the definition of individual transition form factors. Restricting ourselves to the $N_{f}=2$, the parameters $c$, $\sigma$, and $m_{q}=m_{u,d}$ are fitted to the $\rho$-meson mass ($m_{\rho}=2c$), the pion decay constant and mass, respectively, while $g_{5}=2\pi$ follows from a fit of the vector-current correlator to the OPE of QCD.
@@
$\partial_{z}(\frac{e^{-\phi}}{z}\partial_{z}A^{\perp}_{\mu})+\frac{(M_{n}^{A})^{2}}{z}e^{-\phi}A^{\perp}_{\mu}-g_{5}^{2}\frac{v^{2}}{z^{3}}e^{-\phi}A^{\perp}_{\mu}=0.$ (8)
@@
which also leads to a Schrödinger equation, now with a potential (in the chiral limit)
@@
$\tilde{V}_{\mu}(z,q)=\mathcal{J}(z,Q)P_{\mu}^{\;\nu}(q)V^{(0)}_{\nu}(q)+q_{\mu}\frac{q^{\nu}V^{(0)}_{\nu}(q)}{q^{2}},$ (16)
with $V^{(0)}$ being the boundary value and $P^{\mu\nu}(q)=\eta^{\mu\nu}-\frac{q^{\mu}q^{\nu}}{q^{2}}$. The bulk-to-boundary propagator $\mathcal{J}(z,Q)$ with virtuality $Q\equiv\sqrt{-q^{2}}$ obeys the equation
$\partial_{z}\left(\frac{e^{-\phi}}{z}\partial_{z}\mathcal{J}(z,Q)\right)+\frac{q^{2}e^{-\phi}}{z}\mathcal{J}(z,Q)=0,$ (17)
and $\mathcal{J}(0,Q)=1$. The two-point function is given by
$\Pi_{ab}^{\mu\nu}(q)=\frac{1}{g_{5}^{2}}\delta_{ab}P^{\mu\nu}(q)\frac{\partial_{z}\mathcal{J}(z,Q)}{z}\bigg|_{z\to 0}.$ (18)
In order to avoid a pole at $q^{2}=0$ of the form $\Pi^{\mu\nu}_{ab}\sim\delta_{ab}q^{\mu}q^{\nu}/q^{2}$ from massless scalars in $V_{\mu}\propto\partial_{\mu}S$, one needs $z^{-1}{\partial_{z}\mathcal{J}(z,0)}|_{z\to 0}$ to vanish [^2] The familiar divergence ${\partial_{z}\mathcal{J}(\varepsilon,Q)}/{\varepsilon}\sim q^{2}\log(\varepsilon Q)$, which is canceled by holographic renormalization, is not relevant in this discussion as it only appears at nonzero $q^{2}$. . For $a<0$, the most general regular solution of the EOM obeying the UV b.c. at $q^{2}=0$ is
@@
For $a>0$, the second term needs to vanish since it would grow at large $z$, but for $a<0$, any $c_{2}$ seems to be allowed. In order to determine $c_{2}$, we look at small but nonzero values of $q^{2}$ and demand that $\mathcal{J}(z,Q)$ is regular as $q^{2}\rightarrow 0$. It is then immediately seen that when expanding $\mathcal{J}(z,Q)=\mathcal{J}(z,0)+q^{2}\mathcal{J}^{(1)}(z,Q)+$... the perturbation $\mathcal{J}^{(1)}$ grows logarithmically as $z\rightarrow 0$ if $1+c_{2}\int_{0}^{\infty}dz\,z\,e^{\phi}\neq 0$. Demanding regularity means that
@@ III.1 WKB and definition of four regions
However, even at very large $E_{n}$, there will always be regions where $\frac{k^{\prime}}{k^{2}}\ll 1$ is not obeyed, and we need to find different ways to approximate the solution. These regions are the small-$z$ region, where $V_{P,A}(z)\gg E_{n}$ and therefore $\frac{k^{\prime}}{k^{2}}\sim\frac{1}{z^{4}}$ (the size of this region shrinks, as $E_{n}$ grows), and the regions around the points $z_{l},z_{r}$ where $k^{2}(z)=0$. For $V_{P}$, the points $z_{l},z_{r}$ are approximately given by
@@
In appendix A we have spelled out explicitly also the analysis for the axial vector mesons, which differs from the pseudoscalar case in a rather minimal way. The potential $V_{A}$ is slightly different at small $z$, leading also to a mass spectrum $M^{A}_{n}\sim n^{2/3}$, whereas the normalization condition (15) does not involve $E_{n}=(M^{A}_{n})^{2}$.
@@ III.2 Transition form factors, sum rules and hadronic light-by-light tensor
Even without the use of the WKB approximation, one can see signs of this pathological behavior in the g2g-2 contribution by simply computing its value for a sufficiently large number of excited states numerically. The results (with $m_{q}$ and $\sigma$ such that $m_{\pi^{0}}=134.9$ MeV and $f_{\pi}=93.1$ MeV to match [29, 30]) are shown in Table 1 where one can clearly see an increase in the $a_{\mu}$ contributions from excited pseudoscalars with masses above $7.5$ GeV, preventing a convergence of their sum.
@@
The analysis for large $M_{n}^{A}$ is very similar to the pseudoscalar case one, so we will not repeat it in detail here. Going through the steps as in the pseudoscalar case, we find the scaling
@@
The WKB machinery may also be used for hard-wall models to compute transition form factors for large $n$. We give closed formulas for $A_{n}(Q^{2},Q^{2})$ and $A_{n}(Q^{2},0)$ in B and comment on the realization of the Melnikov-Vainshtein constraint. The asymptotic scaling laws of [54] are recovered for $Q\gg\sqrt{E_{n}}=M_{n}^{A}$, but in this limit they do not contribute to the short-distance constraints anymore.
@@ IV VVA correlator and non-commutativity of limits
$\times\bigg(q_{2}^{\alpha}\mathcal{J}^{\prime}(z,Q_{1})\mathcal{J}(z,Q_{1})-q_{1}^{\alpha}\mathcal{J}^{\prime}(z,Q_{2})\mathcal{J}(z,Q_{1})\bigg).$ (66)
$\mathcal{A}^{\perp}(z,q)$ solves the axial equations of motion (8) with $(M_{n}^{A})^{2}$ replaced by $q^{2}$, subject to the boundary condition $\mathcal{A}^{\perp}(0,q)=1$. Similarly $\varphi(z,q)$ together with $\eta(z,q)$ solve the pseudoscalar equations (6) (with $m_{n}^{2}\rightarrow q^{2}$) obeying $\varphi(0,q)=1$, $\eta(0,q)=0$.
@@
$\bar{\mathcal{W}}_{2}^{\mathrm{pole}}(Q^{2})$ $=-\frac{N_{c}}{2\pi^{2}}\text{tr}\left(\mathcal{Q}^{2}\mathcal{Q}_{5}\right)\frac{F_{A}}{Q^{2}+(M^{A})^{2}}\frac{A(Q^{2},0)}{2}$
$=-\frac{N_{c}}{2\pi^{2}}\text{tr}\left(\mathcal{Q}^{2}\mathcal{Q}_{5}\right)\frac{F_{A}}{Q^{2}+(M^{A})^{2}}$
@@
with $F^{A}$ being an axial-vector decay constant and $A(Q^{2},0)$ the singly virtual axial TFF. At large $N_{c}$, the only singularities one would expect in correlation functions and amplitudes (at leading order) are poles, and a dispersive analysis would reconstruct the VVA correlator simply as a sum over all such pole contributions (IV) (except for potential subtractions). According to the discussion in the previous section, the TFF $A_{n}(Q^{2},0)$ scales exponentially, $\sim e^{\frac{c^{2}}{2g_{5}\sigma}\sqrt{E}_{n}}$, and the sum over all pole contributions does not converge. This is seemingly in contradiction to the above analysis, where it was concluded that $\bar{\mathcal{W}}_{2}$ is finite when calculated using the axial bulk-to-boundary propagator. In Appendix E, we trace this back to a lack of uniform convergence of the mode expansion, which does not permit to swap the infinite sum over modes with the indefinite integration over the holographic coordinate $z$.
@@ V Hadronic light-by-light tensor from the 5-dimensional approach
Noting that the Wronskian $\xi_{L}^{\prime}(z^{\prime})\xi_{R}(z^{\prime})-\xi_{L}(z^{\prime})\xi_{R}^{\prime}(z^{\prime})$ is independent of $z^{\prime}$, the symmetry $G_{A}(z,z^{\prime},Q)=G_{A}(z^{\prime},z,Q)$ is apparent. Explicit formulas for $\xi_{L},\,\xi_{R}$ in the WKB approach are given in Appendix D.
@@
By symmetry, we only have to integrate below the diagonal [^8] The other $z,\,z^{\prime}$ dependent factors from the vector bulk-to-boundary propagators are not symmetric in $z,\,z^{\prime}$ for generic $Q_{i}$, but inside the integral only the symmetric part contributes.. Below the diagonal at large $z,z^{\prime}$, the WKB approximation detailed in Appendix D predicts the axial bulk to bulk propagator to scale like
@@ VI.1 Background modifications
for fixed [^11] We have neither determined the power-law prefactor $(M_{n}^{A})^{\lambda}$ nor the $Q_{1},Q_{2}$ dependence, as the exponential already proves the divergence of the HLbL tensor. $Q_{1},Q_{2}$ by similar arguments as in section III.2.
If the second term in (95) contributes at the same order or even dominates, the individual TFFs will also be well defined, but exponential scaling similar to (1) will still occur. Also the axial-vector spectrum will not lie on a linear Regge trajectory (except when $w(z^{2})v(z)^{2}\sim z^{2}$, but then with a different slope than vector mesons). The exponent in the scaling law will in general also not be proportional to $M_{n}^{A}$ anymore, rather if the second term in the potential goes like $z^{\alpha}$, with $\alpha\geq 2$, the exponent will be proportional to $(M_{n}^{A})^{4/\alpha}$. As $\alpha$ gets larger, the factor $e^{(M_{n}^{A})^{4/\alpha}}$ will get smaller at fixed $M_{n}^{A}$, but as $M_{n}^{A}\rightarrow\infty$ it will still dominate any power-law factor and observables such as the light-by-light tensor will diverge. In the case that the $w^{2}(z)v^{2}(z)$ factor behaves exponentially at large $z$, the point $z_{r}$ will scale as $z_{r}\sim\left(\log(E_{n})\right)^{1/b}$, for some $b>0$. In that situation, the scaling of the TFF with $E_{n}$ becomes a power law, and a more sophisticated analysis is needed to decide whether the infinite sums discussed in this paper converge. It seems plausible that there are ranges of $b$ for which convergence can be achieved. Another interesting case is when $w^{2}(z)v^{2}(z)$ has a divergence at some finite $z=z_{0}$. In that case, $z_{r}$ will be equal to $z_{0}$ for all $n$ larger than some critical $n_{c}$. The exponential factor $e^{\frac{c^{2}}{2}z_{r}^{2}}$ is then completely harmless and the spectrum becomes hard-wall like [^12] To have well defined individual TFFs one additionally needs $w(z)^{2}v^{2}(z)\sim az^{\beta}$ for $z>z_{0}$ with $a>0,\beta\geq 2$. . As briefly discussed at the end of the following subsection, the relevant sums all converge for hard-wall models.
@@ VI.2 Scalar-extended CS term
Hard-wall models do not suffer from the above exponential scaling of TFFs since the modes are restricted to a finite region $z\in[0,z_{0}]$. [^14] This is also the case for the Witten-Sakai-Sugimoto model when brought into Schrödinger form. There, however, the geometry is not asymptotically AdS so that one cannot match to the short-distance constraints of QCD. This leads to the scaling behaviors $F(0,0)\sim E_{n}^{-1},f_{n}\sim E_{n}^{-1/4}$ with $E_{n}\sim n^{2}$, which leads to a convergent sum rule (53), and a convergent light-by-light tensor for the pseudoscalars (the axials are analogous). Moreover, the experimental data for the pion TFF (and also for the $a_{1}$) turn out to be reproduced surprisingly well [18, 60, 23] with a standard Chern-Simons term. [^15] In the HW model, a scalar-extended Chern-Simons term with $\alpha^{\prime}=\frac{1}{6}$ helps to improve some features in the $f_{1}$-$f_{1}^{\prime}$ sector, but also leads to an unsatisfactory pion TFF [20].